The information is for the following 3 questions. In a clinical trial, 12 out of 600 patients taking a prescription drug complained of flulike symptoms. Suppose that it is known that 2.5​% of patients taking competing drugs complain of flulike symptoms. Test if the prescription drug cause less flulike symptoms as a side effect at the α=0.02 level of significance.

Estimate the right boundary of the 98% C.I. of the population proportion.

All answers must be rounded to 4 decimal places
Answer #1

Solution :

Given that,

n = 12

x = 600

Point estimate = sample proportion = \hat p = x / n = 12 / 600 = 0.020

1 - \hat p = 0.980

Z\alpha= 2.054

Margin of error = E = Z\alpha * \sqrt ((\hat p * (1 - \hat p )) / n)

= 2.054 * (\sqrt((0.020 * 0.980) / 600)

= 0.0117

Right boundary =  \hat p + E = 0.020 + 0.0117 = 0.0317